### v1

# Beyond PCA: Manifold Dimension Estimation via Local Graph Structure

Abstract
Local principal component analysis (Local PCA) has proven to be an effective tool for estimating the intrinsic dimension of a manifold. More recently, curvature-adjusted PCA (CA-PCA) has improved upon this approach by explicitly accounting for the curvature of the underlying manifold, rather than assuming local flatness. Building on these insights, we propose a general framework for manifold dimension estimation that captures the manifold’s local graph structure by integrating PCA with regression-based techniques. Within this framework, we introduce two representative estimators: quadratic embedding (QE) and total least squares (TLS). Experiments on both synthetic and real-world datasets demonstrate that these methods perform competitively with, and often outperform, state-of-the-art alternatives.
keywords:
manifold hypothesis , intrinsic dimension estimation , nonlinear dimension reduction , ordinary least squares , total least squares
journal: Pattern Recognition
[…]
[unsw] organization=School of Mathematics and Statistics, University of New South Wales, addressline=High St, city=Kensington, postcode=2052, state=NSW, country=Australia
{graphicalabstract}
{highlights}
We introduce a family of estimators for manifold dimension by integrating PCA with regression techniques, enabling effective capture of intrinsic geometric structure.
Extensive experiments on both synthetic and real-world datasets demonstrate that our proposed estimators match or outperform state-of-the-art methods across a wide range of settings.
1 Introduction
Representing high-dimensional data using a significantly smaller number of variables is appealing to both practitioners and theorists. On one hand, many classical methods do not scale well in high-dimensional settings, and working with low-dimensional representations is often more efficient and interpretable. On the other hand, the existence of such representations offers a way to explain the success of modern machine learning techniques, despite the challenges posed by the curse of dimensionality (Keogh and Mueen, 2011).
Traditional dimensionality reduction techniques, such as principal component analysis (PCA, Wold et al. (1987)), typically assume that data points lie in a subspace of the ambient space $\mathbb{R}^{p}$. Hence, there exists a global linear map that provides a low-dimensional representation of the entire dataset.
More recent approaches often rely on the more general manifold hypothesis, which posits that high-dimensional data lie on or near a low-dimensional manifold embedded in the ambient space (Meilă and Zhang, 2024). In this setting, a global coordinate representation may no longer exist; however, the data can still be locally described using only $d$ ($\leq p$) parameters, where $d$ is the dimension of the underlying manifold. We typically refer to $d$ as the intrinsic dimension and $p$ as the ambient dimension.
A natural research problem arising from the manifold hypothesis is to estimate the intrinsic dimension $d$ from data—a task commonly referred to as manifold dimension estimation. Numerous methods have been proposed across a wide range of domains (e.g., Verveer and Duin (1995); Kégl (2002), Yang et al. (2007); Carter et al. (2009); Sricharan et al. (2010), Kalantana and Einbecka (2013); Facco et al. (2017); Lim et al. (2024)), see Bi and Lafaye de Micheaux (2025) for the most recent survey on this topic. Broadly speaking, most manifold dimension estimators are built upon the so-called flatness assumption, which approximates local regions of the underlying manifold as $d$-dimensional linear spaces. This assumption transforms the inherently nonlinear problem into a linear one, enabling the design of numerous estimators.
In the survey, we conducted an extensive set of experiments to benchmark the performance of eight representative manifold dimension estimators. Among them, three emerged as top performers: two are flatness-based—namely, the dimensionality estimator based on angle and norm concentration (DanCo, Ceruti et al. (2014)) and the two-nearest-neighbor estimator (TwoNN, Facco et al. (2017)); the third is the curvature-adjusted PCA estimator (CA-PCA, Gilbert and O’Neill (2025)).
CA-PCA explicitly incorporates curvature information into the local PCA procedure, thereby improving upon the already competitive, though often overlooked, Local PCA estimator (Fukunaga and Olsen, 1971). TwoNN pushes the flatness assumption to its extreme, mitigating curvature effects by relying solely on the two nearest neighbors of each point. DanCo is based on the distribution of distances and angles among data points within neighborhoods, assuming local flatness, and further improves robustness in high-dimensional settings through a Kullback–Leibler (KL) divergence–based comparison framework.
Despite their superior performance relative to other estimators, as discovered in the survey, there remains substantial room for improvement: DanCo is particularly time-consuming; TwoNN tends to underestimate when the intrinsic dimensionality is high; and CA-PCA may overestimate the dimension when the manifold is nonlinearly embedded in an ambient space of much higher dimension.
Motivated by the strengths and limitations of existing estimators, particularly the robustness of PCA-based methods and the improvements of CA-PCA gained by explicitly accounting for curvature—we propose a framework for designing manifold dimension estimators that recovers the local graph structure of the manifold via regression on PCA coordinates extracted from each neighborhood. This approach enables flexible modeling of nontrivial manifold geometry and noise by embedding the problem into a regression setting, rather than attempting to compute geometric quantities directly from sparse data. As a result, it supports more stable and accurate dimension estimation.
Two representative estimators within this framework are described in details: the quadratic embedding estimator (QE) and the total least squares estimator (TLS). Our experiments show that these estimators—despite their simplicity—achieve performance comparable to state-of-the-art methods—and in many cases, especially with small sample sizes or nonlinear embeddings, outperform them.
The remainder of this article is organized as follows. Section 2 reviews fundamental manifold constructions and introduces our design framework. Section 3 presents the proposed estimators, followed by experimental results in Section 4. We conclude in Section 5.
2 Manifold Geometry
In this section, we briefly review several fundamental manifold constructions essential for understanding the methods introduced in this article, before presenting our general design framework for manifold dimension estimation. A more comprehensive treatment of the theoretical foundations is provided in our previous survey. For readers interested in more advanced topics in differential geometry, we refer to Lee (2003) and Lee (2018), which offer detailed expositions of the relevant mathematical background.
2.1 Tangent Space and Local Graph Representation
Throughout this article, we take a manifold $M$ as a $d$-dimensional Riemannian submanifold embedded in an ambient Euclidean space $\mathbb{R}^{p}$, generalizing smooth curves and surfaces to higher dimensions. The intrinsic dimension $d$ then refers to the number of free parameters required to describe the manifold locally. That is, for any point $\bm{x}_{0}\in M$, there exists a neighborhood $U(\bm{x}_{0})\subseteq M$ that can be smoothly mapped to $\mathbb{R}^{d}$ via an injective function $\varphi$, as illustrated in Figure 1(a).
Let $\varphi^{-1}$ denote the inverse of $\varphi$. The tangent space of $M$ at $\bm{x}_{0}$ is defined as the $d$-dimensional vector space (with operations trivially defined with respect to $\bm{x}_{0}$) given by
$T_{\bm{x}_{0}}M=\{\bm{x}_{0}+d\varphi^{-1}\big|_{\varphi(\bm{x}_{0})}(\bm{u}):\bm{u}\in\mathbb{R}^{d}\}.$
Here, $\bm{x}_{0}+d\varphi^{-1}\big|_{\varphi(\bm{x}_{0})}(\bm{u})$ corresponds to the first-order Taylor expansion of $\varphi^{-1}$, making $T_{\bm{x}_{0}}M$ the best linear approximation of the manifold near $\bm{x}_{0}$, as illustrated in Figure 1(b).
The tangent space approximation forms the basis of the widely used flatness assumption in manifold dimension estimation. Suppose we are given a point $\bm{x}_{k}$ along with its $K$-nearest neighbors sampled from $M$. The flatness assumption ignores any geometric or distributional irregularities in the neighborhood and posits the data are uniformly distributed within a $d$-dimensional ball $B_{\bm{x}_{k}}(R)\subseteq T_{\bm{x}_{k}}M$. This local linearization enables the construction of various statistics related to $d$. In this context, “flatness” refers to ignoring curvature at a local scale.
(a) Local representation around $\bm{x}_{0}$.
(b) Tangent space at $\bm{x}_{0}$.
Figure 1: Local coordinate and tangent space.
Furthermore, the neighborhood around $\bm{x}_{0}$ can be reparameterized to better reflect the manifold’s local geometry. By translating $\bm{x}_{0}$ to the origin, we construct a new local coordinate system consisting of $d$ orthonormal vectors spanning the tangent space $T_{\bm{x}_{0}}M$, and pdp-d orthonormal vectors spanning the normal space $T_{\bm{x}_{0}}M^{\perp}$, as shown in Figure 2(a). Standard results from differential geometry (Lee, 2003) then guarantee that any point $\bm{x}$ sufficiently close to $\bm{x}_{0}$ can be expressed in these coordinates as
$\bm{x}^{\prime}=\big(\bm{x}^{\prime}_{1:d},\,g(\bm{x}^{\prime}_{1:d})\big),$ (1)
where $\bm{x}^{\prime}_{a:b}$ denotes the $a$th to $b$th components of $\bm{x}^{\prime}$ (inclusive), $\bm{x}^{\prime}_{0}=0$, and $g\colon\mathbb{R}^{d}\to\mathbb{R}^{p-d}$ is a smooth function satisfying $\nabla g(0)=0$.
(a) Linear Approximation around $\bm{x}_{0}$.
(b) Quadratic Approximation around $\bm{x}_{0}$.
Figure 2: Local Graph Representation.
In this new coordinate system, the neighborhood of $\bm{x}_{0}$ is represented as the graph of a function $g$. The first $d$ components, $\bm{x}^{\prime}_{1:d}$, serve as coordinates in the tangent space, while the remaining pdp-d components are determined by $g$, capturing deviations from the tangent space approximation. The flatness assumption corresponds to the case where $g\doteq 0$, meaning the last pdp-d coordinates vanish—equivalent to retaining only the first-order Taylor expansion of $g$ around the origin. Consequently, the nontrivial local geometry of the manifold are entirely encoded in the higher-order terms of the Taylor expansion of $g$. In particular, the second-order Taylor polynomial of $g$ provides a quadratic approximation, as illustrated in Figure 2(b).
2.2 Linear and Nonlinear Embedding
A particularly challenging scenario for manifold dimension estimators arises when the underlying manifold $M$ is space-filling or nonlinearly embedded in the ambient space (Campadelli et al., 2015). Such manifolds typically exhibit more complex curvature. According to the local graph representation discussed above, since there are $d$ linearly independent directions in the tangent space and pdp-d linearly independent normal directions, a total of d(pd)d(p-d) values are required to fully describe the curvature at any given point. The more complex the curvature, the more data are needed to accurately capture the underlying geometry, making dimension estimation increasingly difficult.
For example, a 3-dimensional sphere linearly embedded in $\mathbb{R}^{6}$ is actually contained within a 4-dimensional subspace, and thus its curvature vanishes in many directions (i.e., $g_{2}$ and $g_{3}$ are zero, where the $g_{j}$ are scalar component functions of $g$). In contrast, consider a 3-dimensional deformed sphere embedded in $\mathbb{R}^{6}$, defined via the mapping
$\varphi^{-1}(\bm{u})=(x_{1},\ldots,x_{6}),$
with
$x_{j}=\left[R+r\cos(2c\pi u_{j})\right]\cos(2\pi u_{j}),\quad x_{j+3}=\left[R+r\cos(2c\pi u_{j})\right]\sin(2\pi u_{j}),$
for $j=1,\ldots,3$, where $R,r,c>0$ are fixed constants. This manifold exhibits nontrivial curvature in all $3\times 3$ tangent and normal directions. Estimating its intrinsic dimension is therefore significantly more difficult due to the pervasive and nonlinear curvature present in all local neighborhoods.
A major weakness of flatness-based estimators such as Local PCA and DanCo is their deteriorated performance on nonlinearly embedded manifolds like the deformed sphere. Given the same number of data points in a neighborhood, the flatness assumption is increasingly violated as the curvature becomes more complex.
2.3 Capturing Local Graph Structure
Let $\mathcal{M}$ denote the set of all manifolds embedded in $\mathbb{R}^{p}$. A natural way to partition $\mathcal{M}$ is by classifying its elements according to their intrinsic dimensions. Alternatively, since any $d$-dimensional manifold can locally be represented as the graph of a smooth function with $d$-dimensional inputs and (pd)(p-d)-dimensional outputs — which we refer to as a $C_{d}^{\infty}$ function — the components of this partition can equivalently be expressed as
$\mathcal{M}_{j}=\left\{\,M\in\mathcal{M}:M\text{ is locally the graph of }C_{j}^{\infty}\text{ functions}\,\right\},$
where $1\leq j\leq p$.
Given a dataset in $\mathbb{R}^{p}$ that satisfies the manifold hypothesis, the underlying manifold must belong to one of $\mathcal{M}_{j}$s. Our approach to manifold dimension estimation is based on recovering the correct local graph structure of this manifold across neighborhoods — that is, identifying the appropriate function class $C_{j}^{\infty}$ from which the local representation functions belong, which can be approached by fitting the data to candidate functions specified for different $\mathcal{M}_{j}$s.
More specifically, let $\{G_{j}\}_{j=1}^{p}$ denote our full candidate model specification, where each $G_{j}$ is a family of functions from $C_{j}^{\infty}$ we choose to model local graph structures of a $j$-dimensional manifold, i.e., $G_{j}$ corresponds to certain classes of $j$-dimensional manifolds. For each candidate dimension $j$, the data is fitted to the corresponding model class $G_{j}$, and a goodness-of-fit score $\sigma_{j}$ is computed to assess the quality of the fit. Ideally, the model is (approximately or exactly) correctly specified only at the true intrinsic dimension $j=d$, which should yield the best fit and, consequently, the highest score. Therefore, a dimension estimator can be constructed as
$\hat{d}=\arg\max_{1\leq j\leq p}\sigma_{j}.$ (2)
Further details about the model fitting procedure and evaluation procedure will be discussed in the next section.
In its most general form, a full model specification may span all $p$ possible partition classes. Within each partition, one may consider multiple families of $C_{j}^{\infty}$ functions, each tailored to model specific local regions of the manifold. The theory guarantees the existence of local graph representation under very mild conditions: it suffices for $g\in C_{j}^{\infty}$ to be a smooth function satisfying $\nabla g(0)=0$. However, this level of generality is typically too broad to support practical estimation or meaningful model comparison.
To make progress, we must constrain the function classes used to fit data within each partition. The choice of these classes should ideally be guided by our prior knowledge of the underlying manifold. In the absence of such information, a natural default is to take each $G_{j}$ as a family of quadratic functions. Specifically, we approximate the output of the graph function $g\colon\mathbb{R}^{j}\to\mathbb{R}^{p-j}$ using a tensor of quadratic forms. Let $\mathcal{Q}\in\mathbb{R}^{(p-j)\times j\times j}$ be a third-order tensor, where each slice $\mathcal{Q}[\ell,:,:]=Q_{\ell}$ is a symmetric matrix encoding the second-order variation of the manifold in the $\ell$-th output direction. Then the functions in $G_{j}$ take the form
$q_{\ell}(\bm{x}_{1:j}^{\prime})=\bm{x}_{1:j}^{\prime\top}Q_{\ell}\bm{x}_{1:j}^{\prime}=\sum_{i=1}^{j}\sum_{k=1}^{j}\mathcal{Q}[\ell,i,k]\,x_{i}^{\prime}x_{k}^{\prime},\quad\ell=1,\ldots,p-j.$ (3)
These quadratic models are not exact in general, but they correspond to the second-order Taylor expansion of the true graph functions, capturing the leading curvature information of the manifold.
We can do better when additional geometric information about the underlying manifold is available. For example, if the manifold is known to be a $d$-dimensional sphere of radius $R$ (with $d$, and possibly $R$, unknown), we can exploit the fact that the local graph function of a sphere has a common structure across different dimensions and local regions:
$g(\bm{x}_{1:j}^{\prime})=\left(R-\sqrt{R^{2}-\|\bm{x}_{1:j}^{\prime}\|^{2}},\,0,\ldots,0\right),$
where the nonzero component captures the curvature of the sphere. Incorporating such prior knowledge can lead to more accurate modeling than generic quadratic approximations.
To summarize, our proposed framework for manifold dimension estimation enables the construction of a family of estimators based on accurately capturing the local graph structure of the underlying manifold. For each candidate dimension, specific local graph models are specified, guided by prior geometric knowledge. These models are then fitted to the data, and their goodness of fit is evaluated to estimate the intrinsic dimension.
Our framework offers several advantages over existing approaches:
  • 1.
    Curvature-aware modeling. By allowing flexible specification of local graph functions—particularly through quadratic forms—we explicitly model the curvature of the underlying manifold. Unlike methods such as CA-PCA, which attempt to infer curvature directly from sparse and potentially noisy local samples, our approach embeds the problem into a regression framework, transforming it into a more stable model-fitting task.
  • 2.
    Model flexibility. The framework permits highly flexible choices of models for different candidate dimensions, and even allows different models to be used for the same dimension across neighborhoods. This makes it possible to incorporate prior geometric knowledge into model specification. On one end, quadratic models provide a natural default; alternatively, one could consider specifying semiparametric or nonparametric models—such as neural networks—to achieve greater flexibility (Park and Sandberg, 1991). However, such models may be impractical due to the typically limited sample size in each neighborhood, or unnecessarily complex for the specific task of estimating the manifold dimension.
  • 3.
    Robustness to noise. The framework naturally accommodates noise modeling, which is critical for real-world datasets. In contrast, many existing manifold dimension estimators are designed under the assumption of clean, noiseless data.
3 Design of New Estimators
The concepts behind QE and TLS follow naturally from our framework. For both estimators, the model specification for candidate dimensions up to p1p-1 is given by the quadratic form in (3). With new coordinates for each neighborhood in hand, the model for dimension $j$ can be fitted by treating the first $j$ coordinates of each data point as inputs and the remaining pjp-j coordinates as outputs. A local PCA procedure is applied to each neighborhood to estimate the new coordinates guaranteed by the theory, which can be reused across all candidate dimensions. The two estimators differ in their choice of fitting and evaluation methods: QE uses ordinary least squares (OLS), while TLS employs total least squares (TLS), which accounts for noise in both input and output directions.
3.1 Regression with PCA Coordinates
Following the above discussion, given a data point $\bm{x}_{k}$ and its $K$-nearest neighbors $\bm{x}_{k}^{1},\ldots,\bm{x}_{k}^{K}$ (ranked by their Euclidean distances from $\bm{x}_{k}$), we first apply PCA to compute the eigendecomposition of the sample covariance matrix of these points. This yields an orthonormal basis $\bm{v}_{1},\ldots,\bm{v}_{p}$ with corresponding eigenvalues $\lambda_{1}\geq\cdots\geq\lambda_{p}$. Taking the local mean $\bar{\bm{x}}_{k}$ as the new origin, each point in the neighborhood can then be expressed in a new coordinate system aligned with the principal directions, denoted by $\bm{x}_{0}^{\prime},\ldots,\bm{x}_{K}^{\prime}$.
(a) Local linear approximation via PCA.
(b) Local linear approximation via $T_{\bm{x}_{k}}M$.
Figure 3: Practical and theoretical local approximations.
By construction, the subspace spanned by $\bm{v}_{1},\ldots,\bm{v}_{d}$ (centered at $\bar{\bm{x}}_{k}$) provides the optimal $d$-dimensional linear approximation of the data points, in the sense that it minimizes the sum of squared distances from the points to the subspace, among all linear spaces of dimension at most $d$. This yields a practical linear approximation of the data points, as shown in Figure 3(a).
On the theoretical side, it is also known that the entire neighborhood around $\bm{x}_{k}$ can be approximated by the tangent space at that point, $T_{\bm{x}_{k}}M$. Under the local graph representation (1), this corresponds to taking the first-order Taylor approximation of the function $g$, which is zero. This approximation is illustrated in Figure 3(b).
Intuitively, and supported by the strong empirical performance of Local PCA—which, at its core, uses the span of principal directions to approximate the tangent space—these two approximations should be close at sufficiently small scales. In other words, the subspace spanned by the first $d$ eigenvectors provides an estimate of the tangent space, up to a small translation and rotation. Indeed, PCA is widely regarded as the most natural method for tangent space estimation; see Lim et al. (2024) for a theoretical discussion of its approximation properties.
For our purposes, the most important consequence is that PCA coordinates serve as good estimates of the theoretical (new) coordinates appearing in the local graph representation, effectively enabling us to fit the models specified for different dimensions. At its core, our estimator seeks to identify the splitting index $d$ at which a deterministic relationship emerges between the first $d$ coordinates and the remaining pdp-d, governed by the local graph function. The correct intrinsic dimension should correspond to a clear improvement in model fit, thereby revealing the true value of $d$.
With the PCA coordinates, recall the quadratic models
$q_{\ell}(\bm{x}_{1:j}^{\prime})=\bm{x}_{1:j}^{\prime\top}Q_{\ell}\bm{x}_{1:j}^{\prime},\quad\ell=1,\ldots,p-j,$ (4)
which, when cross terms among the $x_{j}$’s are included as predictors, can be expressed as a linear model. The model can then be fitted by treating the last pjp-j PCA coordinates as response variables and applying ordinary least squares (OLS).
Our QE estimator adopts this approximation but uses only the first output coordinate (i.e., $x_{k,j+1}$, corresponding to $q_{1}$) as the response. This design has two key advantages: first, PCA coordinates associated with smaller eigenvalues are typically more sensitive to noise; second, it helps avoid spurious relationships between input and output when $j<d$. In contrast, using all of the last pjp-j coordinates as responses introduces the risk of detecting artificial dependencies, which can negatively affect model evaluation.
Our TLS-based estimator (TLS) mirrors the structure of QE, with the key distinction that the regression step is performed using total least squares (TLS) instead of OLS. We also experimented with LASSO (Tibshirani, 1996) and ridge regression (McDonald, 2009) as alternative fitting methods, but found that they yielded inferior performance in this context.
3.2 Filling Details: QE and TLS
Evaluating the model fit is a crucial step within our framework, as (2) suggests, a good fit signals that the intrinsic dimension has likely been reached. For QE, this step is straightforward: classical statistical theory provides the $F$-statistic, specifically designed to compare a target regression model against a trivial one.
Figure 4 shows the distribution of $F$-statistics for $j=1,2,3$, computed from a sample of size $1000$ drawn uniformly from a $d=3$ deformed sphere (with curvature parameter $c=0.01$). A clear jump in the $F$-statistic is observed at $j=3$, accurately identifying the intrinsic dimension.
(a) $j=1$
(b) $j=2$
(c) $j=3$
Figure 4: Distribution of $F$-statistics for different values of $j$.
In practice, it is often more convenient to base the dimension estimate on the $p$-value associated with the $F$-statistic. However, this approach requires care. The validity of the $F$-distribution hinges on the assumption of normally distributed errors, which may not hold in our setting. Two primary sources of deviation from this assumption exist: (i) error introduced by estimating the tangent space via PCA, and (ii) approximation error in modeling the nonlinear map $g$ with a second-order polynomial $q$. Neither of these guarantees normally distributed residuals, and therefore the theoretical $F$-distribution may not strictly apply.
Despite this, the $F$-statistic is well-known for its robustness to such violations (Box and Watson, 1962; Ali and Sharma, 1996). Empirically, we consistently observe a pronounced jump in its value at the true dimension, and applying a fixed $p$-value threshold (e.g., at the 1% level) leads to reliable results across a broad range of datasets and settings.
An alternative goodness-of-fit measure that penalizes model complexity is the adjusted $R^{2}$ score. This statistic is bounded above by 1, with larger values indicating better fit. A negative adjusted $R^{2}$ suggests that the model performs worse than the trivial model.
The complete QE procedure is summarized in Algorithm 1. Beyond the core framework, we incorporate the following refinements:
  • 1.
    Although the first-order Taylor expansion of $g_{1}$ is theoretically zero—so its quadratic approximation (the true model) contains only second-order terms—in practice, our PCA coordinates only approximate the theoretical new coordinates in (1). Therefore, we include both first- and second-order terms in the regression to account for the deviation.
  • 2.
    Since p1p-1 models are fitted in each neighborhood, the procedure can become inefficient when $p\gg d$. In addition to implementation-level optimizations such as parallelization, we limit the number of models per neighborhood to a maximum of $p_{\max}-1$. This threshold is determined by three factors: the ambient dimension $p$; the largest index for which the standard deviation of the new coordinates is sufficiently large (to avoid fitting near-constant responses); and the neighborhood size $K$, as regression becomes unreliable when $K$ is smaller than the number of predictors. Letting $q$ denote the largest nonconstant coordinate index, we define:
    $p_{\max}=\min\left\{p,q,\left\lfloor\sqrt{1+8K}-1\right\rfloor\right\}.$ (5)
  • 3.
    Each neighborhood yields a local estimate $\hat{d}_{k}$, which is the smallest index $j$ such that the $p$-value of the corresponding $F$-statistic falls below 1%. To improve robustness, we further require that all models for indices greater than $j$ also yield $p$-values below the same threshold.
  • 4.
    To compare models across neighborhoods, we record the adjusted $R^{2}$ of the model corresponding to each $\hat{d}_{k}$. This serves as a joint measure of fit and complexity, and should be close to 1 for a well-fitting model. A local estimate is considered valid only if its adjusted $R^{2}$ is positive. The final global estimate $\hat{d}$ is then computed as the weighted average of all valid local estimates $\hat{d}_{k}$, with weights proportional to their adjusted $R^{2}$ scores.
Algorithm 1 QE
1:Input: sample $\{\bm{x}_{k}\}_{k=1}^{n}$; neighborhood size $K$
2:Output: estimated dimension $\hat{d}$
3:for $k=1$ to $n$ do
4:  Find the $K$-nearest neighbors of $\bm{x}_{k}$
5:  Perform PCA on $\bm{x}_{k}$ and its neighbors; express them in new coordinates as $\bm{x}_{0}^{\prime},\dots,\bm{x}_{K}^{\prime}$
6:  Determine $p_{\max}$ according to equation (5)
7:  for $j=1$ to $p_{\max}$ do
8:   Fit a quadratic regression model (via OLS) between the first $j$ coordinates and the $(j+1)$-th
9:   Compute the $F$-statistic $F_{j}$, its $p$-value $p_{j}$, and the adjusted $R^{2}$ score $w_{j}$
10:  end for
11:  Initialize $\hat{d}_{k}=p_{\max}$, $w_{k}^{*}=0$ for $k=1,\cdots,n$
12:  for $j=1$ to $p_{\max}$ do
13:   if $w_{j}>0$ and $p_{i}<1\%$ for all $i\geq j$ then
14:     Set $\hat{d}_{k}=j$, $w_{k}^{*}=w_{j}$
15:     break
16:   end if
17:  end for
18:end for
19:return $\hat{d}=\sum_{k}w_{k}^{*}\hat{d}_{k}\big/\sum_{k}w_{k}^{*}$ (or the unweighted average if all $w_{k}^{*}=0$)
Compared to ordinary least squares (OLS), the statistical properties of total least squares (TLS) are less well understood. In particular, TLS does not naturally support classical inference tools such as the $F$-test or adjusted $R^{2}$, making hypothesis testing and model comparison more challenging. In light of this, we can directly evaluate model fit using the total TLS error instead, denoted by $\sigma_{q_{j}+1}$. This quantity corresponds to the sum of squared orthogonal distances from the data points to the best-fit affine subspace. Here, $q_{j}=j(j-1)/2+2j$ denotes the number of predictors used, accounting for both the linear and quadratic terms derived from the first $j$ PCA coordinates.
As in QE, the quadratic model is expected to become appropriate only when $j\geq d$. While the total TLS error generally decreases as $j$ increases, a more pronounced drop should be observed at the intrinsic dimension $d$. However, this drop is not always sharp or immediately discernible. Figure 5 illustrates this behavior for models with $j=1,\dots,10$, applied to data sampled from a $d=10$ sphere. Although the error reaches its minimum at $j=10$, the decrease from $j=9$ to $j=10$ is relatively modest.
(a) Total TLS error $\sigma_{q_{j}+1}$
(b) Relative drop $\eta_{j}$ in TLS error
Figure 5: Total TLS error and its relative drops as a function of model dimension.
To improve sensitivity in identifying the intrinsic dimension, we consider the relative drop in total error, rather than its absolute value. Specifically, the relative drop from dimension $j$ to $j+1$ is defined as
$\eta_{j+1}=\frac{\sigma_{q_{j}}-\sigma_{q_{j+1}}}{\sigma_{q_{j}}},$ (6)
where $\sigma_{q_{j}}$ denotes the TLS error associated with dimension $j$.
As shown in Figure 5(b), the relative drop $\eta_{j}$ peaks at $j=d$, providing a natural criterion for estimating the local intrinsic dimension $\hat{d}_{k}$. Empirically, we find that this approach is robust for $d\geq 2$. However, it is unable to distinguish the case $d=1$ by design, as there is no $j=0$ model to compare the total error with.
The complete TLS procedure is summarized in Algorithm 2.
Algorithm 2 TLS
1:Input: sample $\{\bm{x}_{k}\}_{k=1}^{n}$; neighborhood size $K$
2:Output: estimated dimension $\hat{d}$
3:for $k=1$ to $n$ do
4:  Find the $K$-nearest neighbors of $\bm{x}_{k}$
5:  Perform PCA on $\bm{x}_{k}$ and its neighbors; express them in new coordinates as $\bm{x}_{0}^{\prime},\dots,\bm{x}_{K}^{\prime}$
6:  Determine $p_{\max}$ according to equation (5)
7:  for $j=1$ to $p_{\max}$ do
8:   Fit a quadratic regression model (via TLS) between the first $j$ coordinates and the $(j+1)$-th
9:   Compute the total error $\sigma_{q_{j}}$
10:  end for
11:  for $j=2$ to $p_{\max}$ do
12:   Compute the relative drop $\eta_{j}$ using equation (6)
13:  end for
14:  Set $\hat{d}_{k}=\arg\max_{j}\{\eta_{j}\}$
15:end for
16:return $\hat{d}=\frac{1}{n}\sum_{k}\hat{d}_{k}$
4 Evaluation
4.1 Setup
In this section, we evaluate the performance of our proposed estimators, QE and TLS, on a diverse collection of synthetic and real-world datasets. For the synthetic datasets, each experiment is repeated $100$ times using randomly generated samples. To assess both bias and variance, we report the mean and standard deviation of the resulting estimates. For comparison, we also include Local PCA, CA-PCA, TwoNN, and DanCo, all of which have been identified as top-performing methods in the existing literature  [^1] Our experiments were conducted using Python. Specifically, we utilized the skdim package implementations for Local PCA, TwoNN, and DanCo, while the remaining estimators, as well as the sampling and testing procedures, were implemented by us. The code is available at: https://github.com/loong-bi/manifold-dimension-estimation. .
With the exception of TwoNN, all evaluated estimators depend on the neighborhood size $K$ as a hyperparameter. The choice of $K$ has a well-documented impact on performance of flatness-based estimators, as it introduces a trade-off: selecting a small $K$ allows for more accurate local linear approximations due to less spread among neighboring points, but may result in insufficient data to reliably capture the underlying geometry. Conversely, a large $K$ can reduce variance but may lead to biased estimates, as the flatness assumption is violated more. For instance, DanCo typically achieves good performance when $K\doteq 10$. The sensitivity of the other estimators to the choice of $K$ is illustrated in Figure 6, based on experiments conducted on datasets of size $1000$ uniformly sampled from a 5-dimensional unit sphere linearly embedded in $\mathbb{R}^{10}$.
Figure 6: Effect of Neighborhood Size $K$.
Several noteworthy patterns can be observed from Figure 6. As expected for an estimator based on the flatness assumption, the estimates produced by Local PCA tend to increase with neighborhood size. This behavior arises because larger neighborhoods capture more of the global geometry, thereby violating the local linear assumption upon which the estimator relies. In contrast, the estimates from CA-PCA and our proposed methods, QE and TLS, remain remarkably stable across the whole range of neighborhood sizes considered. This robustness sets them apart from many existing methods that are sensitive to the choice of $K$ like Local PCA.
Since CA-PCA, QE, and TLS all explicitly incorporate curvature into the estimation process, their robustness to neighborhood size suggests that doing so enables more effective modeling of local geometry across a broad range of scales. This property is particularly advantageous for our proposed estimators, which, by design, require sufficiently large neighborhoods to operate reliably. For instance, as noted in Equation 5, when the intrinsic dimension is $d=10$, a minimum of $15$ data points per neighborhood is necessary to avoid premature truncation by $p_{\max}$. Consequently, it is both essential and beneficial to evaluate our methods over a wide range of neighborhood sizes.
Finally, since the intrinsic dimension is typically unknown in practice, a fair comparison across estimators requires aggregating results over regions of stability in the neighborhood size $K$. Specifically, we adopt the same approach as in our earlier study: an algorithm searches the hyperparameter space for regions where the standard deviation of the estimates is minimized so that the mean estimates remain approximately constant. If no such stable region is identified, we report the average over all candidate hyperparameter values. For further details, see Bi and Lafaye de Micheaux (2025).
4.2 results
Our first set of experiments evaluates the estimators on 18 synthetic datasets, as described in Table 1. These datasets span a wide range of commonly studied manifolds and are identical to those used in our previous survey. Prior evaluations have shown that Local PCA, CA-PCA, DanCo, and TwoNN perform well on this benchmark. The results of our experiments are summarized in Tables 2 to 5.
Tables 2 and 3 report results for noiseless, uniformly sampled datasets with sample sizes $n=500$ and $n=2000$, respectively. Notably, the selected baseline estimators already perform competitively in this setting. Nonetheless, our proposed estimators demonstrate comparable or superior performance in many cases—particularly when the sample size is small, where competing methods tend to exhibit greater bias and variance, especially on nonlinearly embedded manifolds ($M_{41}$ to $M_{43}$).
However, our methods encounter difficulties on trivial manifolds such as hyperballs ($M_{21}$ to $M_{23}$). This is expected, as both QE and TLS are designed to fit local quadratic models, which are ill-suited for trivial manifolds with zero curvature, i.e., the quadratic model (4) misspecifies. In these cases, the estimators revert to their fallback mechanisms, which are less effective.
Tables 4 and 5 present results for the same set of synthetic datasets with added $p$-dimensional Gaussian noise. As expected, the presence of noise leads to a substantial degradation in performance across nearly all estimators. Nevertheless, in many cases, our proposed estimators continue to outperform existing methods, benefiting from the robustness of the underlying regression procedures, which explicitly account for noise. In comparison, most existing estimators base their design on the noiseless setup.
In particular, the performance of our TLS estimator improves significantly with larger sample sizes, indicating that explicitly modeling noise in both the predictors and responses can be beneficial in noisy settings. It is also noteworthy that even the top-performing methods exhibit considerable bias on relatively simple manifolds such as $M_{6}$ to $M_{10}$, underscoring the inherent difficulty of the dimension estimation task in the presence of noise.
One of the main motivations for explicitly incorporating curvature into the design of manifold dimension estimators is the potential for improved performance on nonlinearly embedded manifolds—i.e., manifolds exhibiting nontrivial curvature in all directions. Such cases are particularly challenging for estimators that rely on local flatness assumptions. The deformed spheres ($M_{41}$ to $M_{43}$) serve as representative examples of this category.
To further evaluate this aspect, we conducted experiments on six additional manifolds generated using the skdim library. These manifolds are all nonlinearly embedded in higher-dimensional ambient spaces and are specifically chosen for their increased geometric complexity. Details of these datasets are provided in Table 6, with the corresponding results presented in Tables 7 and 8 for sample sizes $n=500$ and $n=2000$, respectively.
Among these datasets, the top-performing estimators turn out to be TwoNN, QE, and TLS. In contrast, Local PCA and DanCo exhibit weaker performance, which is expected given their reliance on the flatness assumption and lack of explicit curvature modeling. Interestingly, despite accounting for curvature, CA-PCA also performs poorly—particularly on manifolds where the ambient dimension significantly exceeds the intrinsic dimension, often leading to overestimation. This observation is consistent with findings from our previous survey. A possible reason might be its relying—at least partially—on detecting a gap between significant and negligible eigenvalues. When many small but nonzero eigenvalues are present due to curvature or noise, the distinction becomes blurred, often misleading the estimator and degrading performance.
Lastly, we evaluate our estimators on three real-world datasets commonly used in manifold dimension estimation studies. The first is the ISOMAP dataset, which consists of 698 grayscale images ($64\times 64$ pixels) of a face sculpture and is believed to have an intrinsic dimension of approximately 3. The second is a subset of the MNIST dataset, comprising 6,742 grayscale images of the digit ”1” (each of size $28\times 28$), with an estimated intrinsic dimension between 8 and 11. The third is the ISOLET dataset, which contains 6,240 vectors of length 616, representing frequency-based features extracted from audio recordings of spoken English letters, and is believed to have an intrinsic dimension between 16 and 22. The corresponding results are presented in Tables 9, 10, and 11222The results reported for DanCo, shown in brackets, are taken from Ceruti et al. (2014) and Campadelli et al. (2015), which include implementation-specific adjustments designed to stabilize performance. A direct application of DanCo using the skdim library does not reproduce these results; see Bi and Lafaye de Micheaux (2025) for further discussion..
In all three datasets, the ambient dimension substantially exceeds the intrinsic dimension, resulting in significant overestimation by both Local PCA and CA-PCA. While TwoNN provides estimates close to the expected values for the ISOMAP and MNIST datasets, it tends to underestimate the intrinsic dimension for ISOLET, as anticipated given the higher underlying dimensionality. In contrast, the estimates produced by DanCo and QE fall within the expected ranges for all three datasets. Our TLS estimator also yields results that are consistently close to these expected values, further demonstrating its robustness across a range of real-world scenarios.
5 Conclusion
This article introduces a design paradigm for manifold dimension estimators that aims to effectively recover the local graph structure of the manifold by performing regression on PCA coordinates. Two representative estimators—QE and TLS—are described in detail. Both approximate the local graph structure through a quadratic model; QE employs ordinary least squares, while TLS uses total least squares, providing robustness to noise in both predictors and responses. Extensive experiments on synthetic and real-world datasets demonstrate that both methods perform on par with, and often outperform, leading state-of-the-art estimators. Their advantages are especially pronounced in settings with limited sample sizes or nonlinear embeddings in high-dimensional ambient spaces.
Acknowledgements
This research includes computations using the computational cluster Katana supported by Research Technology Services at UNSW Sydney.
This paper is part of Z. Bi’s Ph.D. thesis.
6 Appendix: Tables of Results
Table 1: The 18 manifolds used in Tables 25.
Manifold Description
$M_{11}$ A $5$-dimensional sphere embedded in $\mathbb{R}^{10}$ with $R=1$.
$M_{12}$ A $10$-dimensional sphere embedded in $\mathbb{R}^{20}$ with $R=1$.
$M_{13}$ A $20$-dimensional sphere embedded in $\mathbb{R}^{40}$ with $R=1$.
$M_{21}$ A $5$-dimensional ball embedded in $\mathbb{R}^{10}$ with $R=1$.
$M_{22}$ A $10$-dimensional ball embedded in $\mathbb{R}^{20}$ with $R=1$.
$M_{23}$ A $20$-dimensional ball embedded in $\mathbb{R}^{40}$ with $R=1$.
$M_{31}$ A $5$-dimensional Gaussian density surface embedded in $\mathbb{R}^{10}$ corresponding to $N(0,0.25I_{5})$.
$M_{32}$ A $10$-dimensional Gaussian density surface embedded in $\mathbb{R}^{20}$ corresponding to $N(0,0.25I_{10})$.
$M_{33}$ A $20$-dimensional Gaussian density surface embedded in $\mathbb{R}^{40}$ corresponding to $N(0,0.25I_{20})$.
$M_{41}$ A $3$-dimensional deformed sphere with $c=0.01$.
$M_{42}$ A $3$-dimensional deformed sphere with $c=0.1$.
$M_{43}$ A $3$-dimensional deformed sphere with $c=1$.
$M_{5}$ A $2$-dimensional cylinder embedded in $\mathbb{R}^{4}$.
$M_{6}$ A $1$-dimensional helix embedded in $\mathbb{R}^{3}$.
$M_{7}$ A $2$-dimensional Swiss roll embedded in $\mathbb{R}^{4}$.
$M_{8}$ A $2$-dimensional Mobius strip embedded in $\mathbb{R}^{4}$.
$M_{9}$ A $2$-dimensional torus embedded in $\mathbb{R}^{4}$.
$M_{10}$ A $2$-dimensional hyperbolic surface embedded in $\mathbb{R}^{4}$.
Table 2: Mean dimension estimates ($\pm$ standard deviation) for 18 manifolds (Table 1) with true dimension $d$, based on 100 replicates of $n=500$ uniform samples. Bold indicates the estimate with minimal MSE.
Manifold Local PCA CA-PCA DanCo TwoNN QE TLS $M_{11}(5)$ $5.00$ p m 0.00 $5.03$ p m 0.01 $6.18$ p m 0.26 $4.90$ p m 0.28 $5.00$ p m 0.00 $5.00$ p m 0.00 $M_{12}(10)$ $10.21$ p m 0.02 $10.10$ p m 0.01 $11.16$ p m 0.37 $9.17$ p m 0.53 $10.00$ p m 0.00 $10.00$ p m 0.00 $M_{13}(20)$ $20.68$ p m 0.02 $19.76$ p m 0.02 $20.13$ p m 0.37 $15.90$ p m 0.90 $20.00$ p m 0.00 $19.99$ p m 0.00 $M_{21}(5)$ $4.91$ p m 0.03 $5.00$ p m 0.00 $4.97$ p m 0.13 $4.64$ p m 0.30 $4.72$ p m 0.28 $2.86$ p m 0.04 $M_{22}(10)$ $9.93$ p m 0.03 $9.99$ p m 0.01 $10.03$ p m 0.17 $8.79$ p m 0.51 $9.84$ p m 0.29 $5.19$ p m 0.18 $M_{23}(20)$ $19.93$ p m 0.02 $19.10$ p m 0.04 $20.01$ p m 0.28 $15.53$ p m 0.92 $19.49$ p m 0.44 $18.17$ p m 0.17 $M_{31}(5)$ $5.00$ p m 0.01 $5.00$ p m 0.00 $4.97$ p m 0.13 $4.55$ p m 0.27 $5.00$ p m 0.00 $5.00$ p m 0.00 $M_{32}(10)$ $10.01$ p m 0.02 $10.00$ p m 0.00 $9.55$ p m 0.39 $8.28$ p m 0.51 $10.05$ p m 0.16 $10.00$ p m 0.00 $M_{33}(20)$ $20.07$ p m 0.05 $19.43$ p m 0.04 $19.59$ p m 0.54 $14.75$ p m 0.83 $20.85$ p m 0.22 $20.00$ p m 0.00 $M_{41}(3)$ $3.24$ p m 0.02 $5.83$ p m 0.02 $3.42$ p m 0.12 $3.05$ p m 0.17 $3.00$ p m 0.00 $3.17$ p m 0.03 $M_{42}(3)$ $3.25$ p m 0.02 $5.83$ p m 0.02 $3.44$ p m 0.12 $3.05$ p m 0.17 $3.00$ p m 0.00 $3.13$ p m 0.02 $M_{43}(3)$ $3.44$ p m 0.03 $5.52$ p m 0.05 $3.58$ p m 0.14 $3.06$ p m 0.16 $3.00$ p m 0.00 $3.47$ p m 0.04 $M_{5}(2)$ $1.95$ p m 0.01 $2.00$ p m 0.00 $2.16$ p m 0.03 $1.98$ p m 0.12 $2.00$ p m 0.00 $2.00$ p m 0.00 $M_{6}(1)$ $1.00$ p m 0.00 $2.75$ p m 0.04 $1.01$ p m 0.02 $1.00$ p m 0.07 $1.97$ p m 0.01 $2.00$ p m 0.00 $M_{7}(2)$ $2.78$ p m 0.04 $2.99$ p m 0.01 $2.18$ p m 0.03 $1.97$ p m 0.12 $2.00$ p m 0.00 $2.00$ p m 0.00 $M_{8}(2)$ $1.93$ p m 0.01 $2.00$ p m 0.00 $2.18$ p m 0.03 $1.96$ p m 0.12 $2.00$ p m 0.00 $2.00$ p m 0.00 $M_{9}(2)$ $2.00$ p m 0.01 $3.00$ p m 0.00 $2.16$ p m 0.03 $2.00$ p m 0.13 $2.00$ p m 0.00 $2.00$ p m 0.00 $M_{10}(2)$ $1.95$ p m 0.01 $2.07$ p m 0.01 $2.18$ p m 0.03 $1.99$ p m 0.12 $2.00$ p m 0.00 $2.00$ p m 0.00
Table 3: Mean dimension estimates ($\pm$ standard deviation) for 18 manifolds (Table 1) with true dimension $d$, based on 100 replicates of $n=2000$ uniform samples. Bold indicates the estimate with minimal MSE.
Manifold Local PCA CA-PCA DanCo TwoNN QE TLS $M_{11}(5)$ $5.00$ p m 0.00 $5.00$ p m 0.00 $6.54$ p m 0.17 $4.97$ p m 0.14 $5.00$ p m 0.00 $5.00$ p m 0.00 $M_{12}(10)$ $10.00$ p m 0.00 $10.01$ p m 0.00 $10.92$ p m 0.14 $9.40$ p m 0.24 $10.00$ p m 0.00 $10.00$ p m 0.00 $M_{13}(20)$ $20.17$ p m 0.01 $19.70$ p m 0.01 $19.97$ p m 0.24 $16.74$ p m 0.52 $20.00$ p m 0.00 $19.94$ p m 0.01 $M_{21}(5)$ $4.93$ p m 0.01 $5.00$ p m 0.00 $5.00$ p m 0.06 $4.76$ p m 0.14 $4.31$ p m 0.27 $2.68$ p m 0.02 $M_{22}(10)$ $9.91$ p m 0.02 $9.99$ p m 0.00 $10.00$ p m 0.11 $8.96$ p m 0.22 $9.92$ p m 0.21 $4.13$ p m 0.06 $M_{23}(20)$ $19.79$ p m 0.03 $19.00$ p m 0.03 $19.98$ p m 0.15 $16.17$ p m 0.43 $19.45$ p m 0.39 $18.18$ p m 0.05 $M_{31}(5)$ $5.00$ p m 0.00 $5.00$ p m 0.00 $5.11$ p m 0.08 $4.68$ p m 0.13 $5.00$ p m 0.00 $5.00$ p m 0.00 $M_{32}(10)$ $10.00$ p m 0.00 $10.00$ p m 0.00 $9.71$ p m 0.31 $8.61$ p m 0.25 $10.00$ p m 0.00 $10.00$ p m 0.00 $M_{33}(20)$ $19.99$ p m 0.01 $19.38$ p m 0.02 $20.87$ p m 0.37 $15.41$ p m 0.38 $20.10$ p m 0.19 $20.00$ p m 0.00 $M_{41}(3)$ $3.00$ p m 0.00 $3.11$ p m 0.01 $2.99$ p m 0.05 $3.03$ p m 0.09 $3.00$ p m 0.00 $3.00$ p m 0.00 $M_{42}(3)$ $3.00$ p m 0.00 $3.12$ p m 0.01 $3.01$ p m 0.05 $3.03$ p m 0.09 $3.00$ p m 0.00 $3.00$ p m 0.00 $M_{43}(3)$ $3.01$ p m 0.00 $3.17$ p m 0.01 $3.76$ p m 0.09 $3.03$ p m 0.09 $3.00$ p m 0.00 $3.02$ p m 0.01 $M_{5}(2)$ $1.97$ p m 0.00 $2.00$ p m 0.00 $2.14$ p m 0.02 $1.97$ p m 0.05 $2.00$ p m 0.00 $2.00$ p m 0.00 $M_{6}(1)$ $1.00$ p m 0.00 $1.00$ p m 0.00 $1.00$ p m 0.00 $1.00$ p m 0.03 $1.00$ p m 0.00 $2.00$ p m 0.00 $M_{7}(2)$ $1.98$ p m 0.01 $2.03$ p m 0.00 $2.16$ p m 0.01 $1.97$ p m 0.06 $2.00$ p m 0.00 $2.00$ p m 0.00 $M_{8}(2)$ $1.96$ p m 0.00 $2.00$ p m 0.00 $2.15$ p m 0.01 $1.97$ p m 0.05 $2.00$ p m 0.00 $2.00$ p m 0.00 $M_{9}(2)$ $2.00$ p m 0.00 $2.00$ p m 0.00 $2.15$ p m 0.02 $2.00$ p m 0.06 $2.00$ p m 0.00 $2.00$ p m 0.00 $M_{10}(2)$ $1.98$ p m 0.00 $2.00$ p m 0.00 $2.16$ p m 0.01 $1.98$ p m 0.06 $2.00$ p m 0.00 $2.00$ p m 0.00
Table 4: Mean dimension estimates ($\pm$ standard deviation) for 18 manifolds (Table 1) with true dimension $d$, based on 100 replicates of $n=500$ uniform samples perturbed by Gaussian noise. Bold indicates the estimate with minimal MSE.
Manifold Local PCA CA-PCA DanCo TwoNN QE TLS $M_{11}(5)$ $7.35$ p m 0.32 $7.19$ p m 0.07 $7.49$ p m 0.24 $6.99$ p m 0.35 $8.40$ p m 0.21 $5.55$ p m 0.01 $M_{12}(10)$ $19.06$ p m 0.06 $13.27$ p m 0.08 $14.69$ p m 0.41 $11.98$ p m 0.61 $11.06$ p m 0.15 $11.61$ p m 0.01 $M_{13}(20)$ $35.93$ p m 0.07 $25.73$ p m 0.11 $34.36$ p m 1.01 $20.84$ p m 1.13 $21.05$ p m 0.13 $19.99$ p m 0.00 $M_{21}(5)$ $9.89$ p m 0.03 $8.77$ p m 0.08 $8.22$ p m 0.22 $7.52$ p m 0.43 $5.77$ p m 0.19 $8.15$ p m 0.01 $M_{22}(10)$ $19.50$ p m 0.04 $15.20$ p m 0.12 $14.65$ p m 0.57 $12.25$ p m 0.63 $10.71$ p m 0.12 $11.79$ p m 0.02 $M_{23}(20)$ $36.23$ p m 0.08 $25.89$ p m 0.12 $35.01$ p m 1.25 $21.03$ p m 0.97 $20.38$ p m 0.12 $19.99$ p m 0.00 $M_{31}(5)$ $5.06$ p m 0.03 $5.04$ p m 0.02 $6.02$ p m 0.15 $5.66$ p m 0.33 $5.74$ p m 0.15 $6.06$ p m 0.01 $M_{32}(10)$ $10.22$ p m 0.08 $10.00$ p m 0.00 $10.16$ p m 0.28 $8.93$ p m 0.50 $11.04$ p m 0.11 $10.64$ p m 0.00 $M_{33}(20)$ $21.15$ p m 0.24 $19.55$ p m 0.03 $20.42$ p m 0.57 $15.28$ p m 0.76 $20.10$ p m 0.19 $19.99$ p m 0.00 $M_{41}(3)$ $3.44$ p m 0.03 $5.90$ p m 0.01 $4.04$ p m 0.14 $3.47$ p m 0.18 $3.00$ p m 0.00 $3.06$ p m 0.01 $M_{42}(3)$ $3.46$ p m 0.03 $5.90$ p m 0.01 $4.07$ p m 0.14 $3.48$ p m 0.18 $3.01$ p m 0.01 $3.03$ p m 0.01 $M_{43}(3)$ $3.89$ p m 0.05 $5.79$ p m 0.03 $4.73$ p m 0.19 $3.88$ p m 0.19 $3.71$ p m 0.10 $3.58$ p m 0.06 $M_{5}(2)$ $3.52$ p m 0.10 $3.93$ p m 0.01 $4.00$ p m 0.00 $3.85$ p m 0.21 $3.50$ p m 0.36 $2.79$ p m 0.02 $M_{6}(1)$ $1.73$ p m 0.04 $2.97$ p m 0.01 $2.94$ p m 0.09 $2.77$ p m 0.15 $1.96$ p m 0.01 $2.00$ p m 0.00 $M_{7}(2)$ $2.78$ p m 0.04 $2.99$ p m 0.01 $2.27$ p m 0.04 $2.18$ p m 0.11 $2.49$ p m 0.45 $2.45$ p m 0.02 $M_{8}(2)$ $3.20$ p m 0.10 $3.94$ p m 0.02 $4.00$ p m 0.02 $3.75$ p m 0.20 $3.04$ p m 0.38 $2.81$ p m 0.02 $M_{9}(2)$ $2.10$ p m 0.01 $3.00$ p m 0.00 $2.89$ p m 0.13 $2.98$ p m 0.15 $2.94$ p m 0.67 $2.74$ p m 0.02 $M_{10}(2)$ $2.15$ p m 0.04 $3.19$ p m 0.06 $3.86$ p m 0.16 $3.29$ p m 0.18 $3.14$ p m 0.39 $2.95$ p m 0.01
Table 5: Mean dimension estimates ($\pm$ standard deviation) for 18 manifolds (Table 1) with true dimension $d$, based on 100 replicates of $n=2000$ uniform samples perturbed by Gaussian noise. Bold indicates the estimate with minimal MSE.
Manifold Local PCA CA-PCA DanCo TwoNN QE TLS $M_{11}(5)$ $9.78$ p m 0.04 $8.40$ p m 0.03 $10.00$ p m 0.02 $7.77$ p m 0.20 $7.33$ p m 0.01 $5.52$ p m 0.01 $M_{12}(10)$ $19.60$ p m 0.02 $14.56$ p m 0.05 $14.89$ p m 0.21 $13.03$ p m 0.32 $13.00$ p m 0.01 $10.60$ p m 0.01 $M_{13}(20)$ $35.38$ p m 0.04 $26.17$ p m 0.05 $36.49$ p m 0.90 $22.68$ p m 0.64 $21.00$ p m 0.00 $21.00$ p m 0.00 $M_{21}(5)$ $9.99$ p m 0.00 $9.61$ p m 0.02 $10.00$ p m 0.00 $8.25$ p m 0.24 $5.67$ p m 0.03 $8.09$ p m 0.03 $M_{22}(10)$ $19.83$ p m 0.01 $16.21$ p m 0.06 $15.02$ p m 0.13 $13.33$ p m 0.30 $11.66$ p m 0.01 $12.75$ p m 0.01 $M_{23}(20)$ $35.68$ p m 0.04 $26.38$ p m 0.06 $37.51$ p m 0.96 $22.81$ p m 0.58 $20.00$ p m 0.00 $19.99$ p m 0.00 $M_{31}(5)$ $5.08$ p m 0.02 $5.87$ p m 0.05 $6.97$ p m 0.09 $6.34$ p m 0.17 $6.33$ p m 0.01 $6.16$ p m 0.01 $M_{32}(10)$ $10.09$ p m 0.02 $10.00$ p m 0.00 $10.49$ p m 0.33 $9.47$ p m 0.27 $11.66$ p m 0.01 $10.63$ p m 0.00 $M_{33}(20)$ $20.27$ p m 0.04 $19.52$ p m 0.02 $22.82$ p m 0.62 $16.13$ p m 0.37 $20.33$ p m 0.00 $19.99$ p m 0.00 $M_{41}(3)$ $3.03$ p m 0.00 $3.85$ p m 0.02 $4.46$ p m 0.12 $3.93$ p m 0.10 $3.00$ p m 0.00 $3.00$ p m 0.00 $M_{42}(3)$ $3.04$ p m 0.01 $3.89$ p m 0.02 $4.52$ p m 0.13 $3.96$ p m 0.10 $3.00$ p m 0.00 $3.00$ p m 0.00 $M_{43}(3)$ $3.53$ p m 0.04 $4.55$ p m 0.03 $5.04$ p m 0.19 $4.47$ p m 0.14 $3.17$ p m 0.03 $3.68$ p m 0.03 $M_{5}(2)$ $3.98$ p m 0.00 $4.00$ p m 0.00 $4.00$ p m 0.00 $4.00$ p m 0.12 $3.07$ p m 0.16 $2.77$ p m 0.01 $M_{6}(1)$ $2.54$ p m 0.01 $2.85$ p m 0.01 $3.00$ p m 0.00 $3.01$ p m 0.08 $2.05$ p m 0.14 $2.00$ p m 0.00 $M_{7}(2)$ $1.99$ p m 0.01 $2.13$ p m 0.01 $2.39$ p m 0.02 $2.50$ p m 0.07 $2.52$ p m 0.15 $2.67$ p m 0.01 $M_{8}(2)$ $3.96$ p m 0.01 $4.00$ p m 0.00 $4.00$ p m 0.00 $3.97$ p m 0.11 $3.14$ p m 0.23 $2.77$ p m 0.01 $M_{9}(2)$ $2.17$ p m 0.02 $3.52$ p m 0.02 $3.98$ p m 0.07 $3.54$ p m 0.10 $2.66$ p m 0.16 $2.95$ p m 0.01 $M_{10}(2)$ $3.07$ p m 0.04 $3.87$ p m 0.01 $4.00$ p m 0.00 $3.76$ p m 0.10 $3.04$ p m 0.14 $2.87$ p m 0.01
Table 6: The 6 nonlinearly embedded manifolds used in Tables 78.
Manifold Description
$M_{\text{NL}1}$ A $6$-dimensional paraboloid nonlinearly embedded in $\mathbb{R}^{21}$.
$M_{\text{NL}2}$ A $4$-dimensional manifold (concentrated figure) nonlinearly embedded in $\mathbb{R}^{6}$.
$M_{\text{NL}3}$ A $4$-dimensional manifold nonlinearly embedded in $\mathbb{R}^{6}$.
$M_{\text{NL}4}$ A $10$-dimensional manifold nonlinearly embedded in $\mathbb{R}^{40}$.
$M_{\text{NL}5}$ A $12$-dimensional manifold nonlinearly embedded in $\mathbb{R}^{72}$.
$M_{\text{NL}6}$ An $18$-dimensional manifold nonlinearly embedded in $\mathbb{R}^{72}$.
Table 7: Mean dimension estimates ($\pm$ standard deviation) for 6 nonlinearly embedded manifolds (Table 6) with true dimension $d$, based on 100 replicates of $n=500$ samples. Bold indicates the estimate with minimal MSE.
Manifold Local PCA CA-PCA DanCo TwoNN QE TLS $M_{\text{NL}1}(6)$ $7.26$ p m 0.21 $5.68$ p m 0.05 $5.48$ p m 0.16 $3.61$ p m 0.06 $5.99$ p m 0.01 $6.36$ p m 0.03 $M_{\text{NL}2}(4)$ $4.64$ p m 0.09 $5.16$ p m 0.07 $4.90$ p m 0.20 $3.67$ p m 0.05 $4.66$ p m 0.05 $4.56$ p m 0.05 $M_{\text{NL}3}(4)$ $7.65$ p m 0.05 $7.44$ p m 0.06 $6.00$ p m 0.28 $4.75$ p m 0.04 $5.26$ p m 0.10 $6.93$ p m 0.01 $M_{\text{NL}4}(10)$ $24.17$ p m 0.84 $9.08$ p m 0.11 $6.36$ p m 0.19 $4.78$ p m 0.10 $9.08$ p m 0.15 $10.96$ p m 0.02 $M_{\text{NL}5}(12)$ $25.25$ p m 0.36 $21.06$ p m 0.08 $18.71$ p m 0.79 $11.52$ p m 0.12 $11.98$ p m 0.02 $12.92$ p m 0.01 $M_{\text{NL}6}(18)$ $25.52$ p m 0.25 $18.04$ p m 0.03 $18.28$ p m 0.74 $11.25$ p m 0.09 $18.00$ p m 0.00 $18.00$ p m 0.00
Table 8: Mean dimension estimates ($\pm$ standard deviation) for 6 nonlinearly embedded manifolds (Table 6) with true dimension $d$, based on 100 replicates of $n=2000$ samples. Bold indicates the estimate with minimal MSE.
Manifold Local PCA CA-PCA DanCo TwoNN QE TLS $M_{\text{NL}1}(6)$ $6.34$ p m 0.06 $5.91$ p m 0.02 $6.61$ p m 0.15 $4.63$ p m 0.03 $5.99$ p m 0.00 $6.36$ p m 0.02 $M_{\text{NL}2}(4)$ $4.44$ p m 0.04 $4.98$ p m 0.03 $5.09$ p m 0.20 $3.92$ p m 0.03 $4.59$ p m 0.02 $4.72$ p m 0.02 $M_{\text{NL}3}(4)$ $4.42$ p m 0.02 $6.28$ p m 0.03 $5.53$ p m 0.18 $4.42$ p m 0.02 $4.16$ p m 0.02 $6.87$ p m 0.01 $M_{\text{NL}4}(10)$ $16.07$ p m 0.32 $9.16$ p m 0.04 $7.00$ p m 0.05 $5.42$ p m 0.04 $9.34$ p m 0.05 $10.92$ p m 0.01 $M_{\text{NL}5}(12)$ $23.92$ p m 0.04 $20.33$ p m 0.05 $17.36$ p m 0.31 $12.77$ p m 0.06 $10.98$ p m 0.01 $12.92$ p m 0.00 $M_{\text{NL}6}(18)$ $22.31$ p m 0.06 $17.89$ p m 0.01 $20.73$ p m 0.53 $12.73$ p m 0.05 $17.00$ p m 0.01 $18.00$ p m 0.00
Table 9: Dimension Estimate on Real-world Datasets: ISOMAP
Local PCA CA-PCA DanCo TwoNN QE TLS
$10.05$ $25.71$ ($4.00$) $3.49$ $3.07$ $3.44$
Table 10: Dimension Estimate on Real-world Datasets: MNIST
Local PCA CA-PCA DanCo TwoNN QE TLS
$15.41$ $84.66$ ($9.98$) $12.98$ $9.23$ $6.73$
Table 11: Dimension Estimate on Real-world Datasets: ISOLET
Local PCA CA-PCA DanCo TwoNN QE TLS
$34.44$ $54.22$ ($19.00$) $9.11$ $17.51$ $25.42$